

********************************************************************
UARS WEEKLY STATUS REPORT - 05 OCT 2001
********************************************************************
***********
SOLSTICE
***********
After resuming normal operations on Monday, October 1st, SOLSTICE continues
taking solar data on all available orbital tracking periods and stellar
calibration observations whenever possible. We continue to gather enough
data to meet all the SOLSTICE science objectives during this period of
reduced tape recorder operations.
***********
SUSIM
***********
When on, SUSIM completed daily mid resolution and MgII solar scans during
periods of predicted UARS telemetry.
We are testing the PACOR-A level 0 SUSIM science files.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing recieved.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG attended the UARS Monthly Status Review Meeting
held on 4 October 2001.
The MPG will continue to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems. In fact, Command Mangement
System processing requires Daily Activity Plans each day
for all operational instruments to maintain continuity
and history of the uplinked command procedures.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 22 September 2001 (Orbit 54841, GMT Day 265) through 28
September 2001 (Orbit 54945, GMT Day 271).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 265/0046 54841 Power ON
WINDII 267/2305 54885 Power OFF
SOLSTICE 267/2305 54885 Power OFF(low pwr)
SUSIM 267/2305 54885 Power OFF
The beta angle ranged between the angles of 40.1 degrees to 53.4 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 54 degrees on 30 September 2001 (DOY 273).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.6 to 1.0 degrees C. Additional battery performance data for each
battery on 21 September 2001 and 26 September 2001 are provided in Appendix D.
The clock error ranged from Š12.3 to 11.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
„ OUT-OF-LIMITS = 00
„ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
„ FOT = 12
„ WINDII = 02
„ SUSIM = 01
„ SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
„ FOT = 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/243 - 2287.499091
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/270 show the observatory data loss to be 303 hours, 12
minutes, and 55.024 seconds. This shows a data loss of 1 minute and 42.320
seconds since the last report period. This is a 0.39 percent data loss, which
equals a 99.61 percent data capture for the mission.
The increase of 1 minute and 42.320 seconds occurred over one (1) days:
· 1 minute and 42.320 seconds on Day 270 due to a Data Hit
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT has begun the process of implementing RaytheonÕs Eclipse ground system
and continues to work on the TAC replacement system. These new systems will
provide enhanced capability and make up for the loss of some UARS facilities,
which have been shut down due to budgetary reasons. This will be a huge effort
and must be completed as soon as possible in order to provide a low cost method
for flying UARS. The FOT will need assistance from all entities involved to
insure the success of this operation. The estimated time for completion is
April 2002.
The FOT has also begun the determination of the overall requirements necessary
to meet the needs of the new ground system as a whole. The requirements will
be examined to determine which entity can better meet the need. The
requirements are also being examined in parallel by Raytheon, so that they too
can determine those which their product, Eclipse, can handle best.
MEETINGS
The UARS FOT attended a PACOR-A ORR on Wednesday, September 26, 2001.
The UARS FOT attended a UARS Requirements Review on Wednesday, September 26,
2001.
The UARS FOT attended a UARS SERB on Wednesday, September 26, 2001.
The UARS FOT attended a UARS Ground System Meeting on Wednesday, September 26,
2001.
The UARS FOT attended a TAC-Eclipse Discussion on Thursday, September 27, 2001.
The UARS FOT attended UARS Technical Team Meeting on Thursday, September 27,
2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 October 2001 (DOY 290). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 September 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
54841 265/0046 01-585 WINDII Power ON
54846 265/0747 01-584 NBTR NOOPS
54852 265/1804 01-586 NBTR NOOPS D266
54860 266/0603 01-586 NBTR NOOPS
54862 266/1024 01-587 OBC Clk Adj fr 0.00 to 20.6 msec/day
54867 266/1804 01-588 NBTR NOOPS D267
54875 267/0605 01-588 NBTR NOOPS
54882 267/1805 01-588 NBTR NOOPS D267
54882 267/1805 01-589 NBTR NOOPS D268
54885 267/2305 01-592 WINDII Power OFF
54885 267/2305 01-593 SOLSTICE Power OFF(low pwr)
54885 267/2305 01-594 SUSIM Power OFF
54887 268/0229 01-595 OBC Clk Adj fr 20.6 to 0.00 msec/day
54891 268/0747 01-589 NBTR NOOPS
54909 269/1246 01-596 NBTR PLAYBACK
54924 270/1215 01-597 OBC Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 September 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 September 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
54862 266/1025 Clk Adj fr 0.00 to 20.6 msec/day
54887 268/0230 Clk Adj fr 20.6 to 0.00 msec/day
54924 270/1218 Clk Adj fr 0.00 to 20.6 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 September 2001
21 September 2001 - GMT Day 264 Beta = 36.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 11.8 0.000 0.0
2 +00.0/-190.4 +5.47/+4.51 +28.0/-14.4 25.1 1.030 19.2
3 +22.4/-44.8 +4.51/+3.23 +27.2/-14.8 25.1 1.030 18.8
26 September 2001 - GMT Day 269 Beta = 53.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.36/-3.99 +0.0/+0.0 11.8 0.000 0.0
2 +00.0/-22.4 +3.23/+2.60 +13.6/-13.2 28.3 1.020 17.8
3 +11.2/-22.4 +2.60/+1.66 +13.6/-12.8 28.3 1.010 16.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
___________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 September 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-054 FOT 54841 265/0046 14170 Generic Late Acquisition
01-054 FOT 54878 267/1106 14170 Generic Late Acquisition
01-058 FOT 54883 267/1904 24053 FWD Failover
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-058 FOT 54883 267/1904 24053 FWD Failover
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-054 FOT see list Sept. 01 14170 Generic Late Acquisition
SEVNTFW01/39